# To the Noise and Back: Diffusion for Shared Autonomy

### v3

@@ B-B Block Pushing
[$p_{\text{noisy}}=0.2$]Refer to caption [$p_{\text{noisy}}=0.3$]Refer to caption [$p_{\text{noisy}}=0.4$]Refer to caption [$p_{\text{noisy}}=0.5$]Refer to caption  [$p_{\text{noisy}}=0.6$]Refer to caption [$p_{\text{noisy}}=0.7$]Refer to caption [$p_{\text{noisy}}=0.8$]Refer to caption [$p_{\text{noisy}}=0.9$]Refer to caption
@@
[$p_{\text{laggy}}=0.2$]Refer to caption [$p_{\text{laggy}}=0.3$]Refer to caption [$p_{\text{laggy}}=0.4$]Refer to caption [$p_{\text{laggy}}=0.5$]Refer to caption  [$p_{\text{laggy}}=0.6$]Refer to caption [$p_{\text{laggy}}=0.7$]Refer to caption [$p_{\text{laggy}}=0.8$]Refer to caption [$p_{\text{laggy}}=0.9$]Refer to caption
@@
Pilot
w/o Copilot ($\gamma=0.0$)
w/ Copilot (ours) ($\gamma=0.2$)
w/ Copilot ($\gamma=1.0$)
w/o Copilot ($\gamma=0.0$)
w/ Copilot (ours) ($\gamma=0.2$)
w/ Copilot ($\gamma=1.0$)
Noisy
$62.33\pm 2.49$
$\bm{75.67\pm 2.05}$
$52.67\pm 7.41$
$\hphantom{0}37.33\pm 2.05$
$\bm{20.00\pm 3.27}$
$5.67\pm 1.25$
Laggy
$42.00\pm 4.90$
$\bm{74.33\pm 2.87}$
$53.33\pm 4.19$
$\hphantom{0}58.00\pm 4.90$
$\bm{21.33\pm 4.50}$
$5.67\pm 1.70$
Zero
$0\hphantom{0}0.00\pm 0$ 0.00
$40.67\pm 2.62$
$52.33\pm 0.47$
$100.00\pm 0.00$
$\hphantom{0}6.67\pm 1.70$
$3.67\pm 2.36$
Random
$0\hphantom{0}0.00\pm 0$ 0.00
$16.33\pm 3.09$
$52.00\pm 3.74$
$100.00\pm 0.00$
$68.00\pm 1.63$
$4.33\pm 2.05$
Expert
$99.00\pm 0.82$
$94.67\pm 2.05$
$51.00\pm 2.45$
$\hphantom{00}1.00\pm 0.82$
$\hphantom{0}5.33\pm 2.05$
$5.67\pm 2.49$

### v4

@@ B-B Block Pushing
$p_{\text{noisy}}=0.2$]Refer to caption $p_{\text{noisy}}=0.3$]Refer to caption $p_{\text{noisy}}=0.4$]Refer to caption $p_{\text{noisy}}=0.5$]Refer to caption  $p_{\text{noisy}}=0.6$]Refer to caption $p_{\text{noisy}}=0.7$]Refer to caption $p_{\text{noisy}}=0.8$]Refer to caption $p_{\text{noisy}}=0.9$]Refer to caption
@@
$p_{\text{laggy}}=0.2$]Refer to caption $p_{\text{laggy}}=0.3$]Refer to caption $p_{\text{laggy}}=0.4$]Refer to caption $p_{\text{laggy}}=0.5$]Refer to caption  $p_{\text{laggy}}=0.6$]Refer to caption $p_{\text{laggy}}=0.7$]Refer to caption $p_{\text{laggy}}=0.8$]Refer to caption $p_{\text{laggy}}=0.9$]Refer to caption
@@
Pilot w/o Copilot ($\gamma=0.0$) w/ Copilot (ours) ($\gamma=0.2$) w/ Copilot ($\gamma=1.0$) w/o Copilot ($\gamma=0.0$) w/ Copilot (ours) ($\gamma=0.2$) w/ Copilot ($\gamma=1.0$)
Noisy $62.33\pm 2.49$ $\bm{75.67\pm 2.05}$ $52.67\pm 7.41$ $\hphantom{0}37.33\pm 2.05$ $\bm{20.00\pm 3.27}$ $5.67\pm 1.25$
Laggy $42.00\pm 4.90$ $\bm{74.33\pm 2.87}$ $53.33\pm 4.19$ $\hphantom{0}58.00\pm 4.90$ $\bm{21.33\pm 4.50}$ $5.67\pm 1.70$
Zero $\hphantom{0}0.00\pm 0.00$ $40.67\pm 2.62$ $52.33\pm 0.47$ $100.00\pm 0.00$ $\hphantom{0}6.67\pm 1.70$ $3.67\pm 2.36$
Random $\hphantom{0}0.00\pm 0.00$ $16.33\pm 3.09$ $52.00\pm 3.74$ $100.00\pm 0.00$ $68.00\pm 1.63$ $4.33\pm 2.05$
Expert $99.00\pm 0.82$ $94.67\pm 2.05$ $51.00\pm 2.45$ $\hphantom{00}1.00\pm 0.82$ $\hphantom{0}5.33\pm 2.05$ $5.67\pm 2.49$